OpenStudy (anonymous):

can any one help me with this one step equation x-5and5 over 8=3 and 1over 2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

(x-5)+5/8 = 3+(1/2)????

OpenStudy (anonymous):

fractions does that help

OpenStudy (anonymous):

(x-5 5/8)=3 1/2???????????????/

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so, 5 5/8 is a mixed number?

OpenStudy (anonymous):

sooo add 5 5/8 to each side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=39/8

OpenStudy (anonymous):

but dont we have to make it in to a mixed number

OpenStudy (anonymous):

you can have improper fraction first/.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

wait what is the final answer

OpenStudy (anonymous):

x=39/8

OpenStudy (anonymous):

but you have to make it a mixed number after though .right....

OpenStudy (anonymous):

ah, yeah the answer is 4 and 7/8

OpenStudy (anonymous):

4 7/8

OpenStudy (anonymous):

thank you again you have helped me a lot

OpenStudy (anonymous):

it is my pleasure to help.